OpenStudy (gigi1103):

Which has a greater velocity? A ball rolling down a 3.4m hill in 8.5s or a fish swimming upstream, covering 6.2m in 0.6 minutes? (velocity = change in distance / change in time)

OpenStudy (gigi1103):

@5sos

OpenStudy (5sos):

They would have the same velocity.

OpenStudy (gigi1103):

for real t thought the fish had a higher one

OpenStudy (mangotangochick):

v = d/t Ball: 3.4/8.5 =0.4ms^-1 v=d/t Fishy: 6.2/(0.6x60) Because to find the number of seconds you do the decimal times the number of seconds in a minute =0.172ms^1

OpenStudy (5sos):

Nope they cant compare because they are 2 different things.

OpenStudy (gigi1103):

wait so there is no answer?

OpenStudy (mangotangochick):

You would probably have to add a sentence to each of them saying where the positive direction is (i.e. for the fish it would be upstream, but for the ball it would be down the hill)

OpenStudy (gigi1103):

my answer choices are They have the same velocity the fish the ball they can't compare because they have different time units

OpenStudy (5sos):

You are dealing wiht 2 different "environments" you could say. It would be D!

OpenStudy (gigi1103):

thank you!!

OpenStudy (5sos):

OFC!!<3